OpenStudy (anonymous):

585 is 150% of what number?

OpenStudy (saifoo.khan):

585 : 150 x : 100

OpenStudy (saifoo.khan):

Cross multiply.

OpenStudy (amistre64):

divide it by 150 and multiply by 100

OpenStudy (anonymous):

390

OpenStudy (anonymous):

Cross mult? I dont understand? :/

OpenStudy (saifoo.khan):

\[x(150) = 585*100\]

OpenStudy (amistre64):

585 = 150% /150 /150 ------------ 3.9 = 1% *100 *100 ------------ 390 = 100%

OpenStudy (saifoo.khan):

Solve for x

OpenStudy (saifoo.khan):

Agree with e-shan.

OpenStudy (amistre64):

if you insist on jking around like that i will have to suspend you. Openstudy is free and any confusion otherwise will have to be snipped in the bud.

OpenStudy (amistre64):

the burden of being a moderator :/

OpenStudy (anonymous):

oh sorry ! i was gonna type in the answer but then you all did it first sorry that won't happen again

OpenStudy (anonymous):

Okay! I got 3,500? :)

OpenStudy (anonymous):

3500!!! do it again

OpenStudy (anonymous):

I did it wrong??? How? Okay h.o!

OpenStudy (anonymous):

0.9t=3150 right?

OpenStudy (saifoo.khan):

wrong^

OpenStudy (anonymous):

585 is 150 % of a number x that means 585 is greater than x

OpenStudy (anonymous):

the original problem was " 585 is 150% of what number?

OpenStudy (anonymous):

Is it 390?

OpenStudy (anonymous):

yea i think so

OpenStudy (anonymous):

Okay thankyou :))